# KAN or MLP? Point Cloud Shows the Way Forward

### v2

@@ Abstract
Multi-Layer Perceptron (MLP) have become one of the fundamental architectural component in point cloud analysis due to its effective feature learning mechanism. However, when processing complex geometric structures in point clouds, MLP’ fixed activation functions struggle to efficiently capture local geometric features, while suffering from poor parameter efficiency and high model redundancy. In this paper, we propose PointKAN, which applies Kolmogorov-Arnold Network (KAN) to point cloud analysis tasks to investigate their efficacy in hierarchical feature representation. First, we introduce a Geometric Affine Module (GAM) to transform local features, improving the model’s robustness to geometric variations. Next, in the Local Feature Processing (LFP), a parallel structure extracts both group-level features and global context, providing a rich representation of both fine details and overall structure. Finally, these features are combined and processed in the Global Feature Processing (GFP). By repeating these operations, the receptive field gradually expands, enabling the model to capture complete geometric information of the point cloud. To overcome the high parameter counts and computational inefficiency of standard KAN, we develop Efficient-KAN in the PointKAN-elite variant, which significantly reduces parameters while maintaining accuracy. Experimental results demonstrate that PointKAN outperforms PointMLP on benchmark datasets such as ModelNet40, ScanObjectNN, and ShapeNetPart, with particularly strong performance in Few-shot Learning task. Additionally, PointKAN achieves substantial reductions in parameter counts and computational complexity (FLOPs). This work highlights the potential of KAN-based architectures in 3D vision and opens new avenues for research in point cloud understanding.
Figure 1: PointKAN (including PointKAN-elite) comprehensively outperforms counterparts. (a) Surpasses MLP and CNN baselines across datasets; (b) In Few-shot Learning, exceeds Point-BERT’s pre-trained performance when trained from scratch; (c)-(d) Substantially reduces parameters versus PointMLP, with PointKAN-elite significantly curbing FLOPs growth as Patch Sequence Length increases.
@@ 1 Introduction
Point cloud analysis is vital in computer vision and 3D processing, with applications spanning autonomous driving  (Li, Luo, and Yang 2023) and robotics (Yan et al. 2020; Rusu and Cousins 2011). Unlike structured 2D images, point clouds pose challenges due to their irregularity and sparsity. Recent deep learning advances (Zhang et al. 2022; Pang et al. 2022; He et al. 2024; Liang et al. 2024; Han et al. 2024) build upon pioneering MLP-based architectures like PointNet/PointNet++  (Qi et al. 2017a, b). Subsequent research, including PointNeXt  (Qian et al. 2022), DGCNN  (Wang et al. 2019), and Point Transformer  (Zhao et al. 2021), employs CNN, GCN, and attention to capture local geometric features.
@@
Notably, PointMLP (Ma et al. 2022) utilizes a pure MLP architecture for point cloud analysis. Its core design philosophy eschews complex convolutions or graph operations, instead relying on stacked residual MLP blocks for feature learning. While achieving competitive performance in classification and segmentation, PointMLP inherits inherent MLP limitations. Rooted in the Universal Approximation Theorem (Cybenko 1989), MLPs employ linear weights and fixed activations. Consequently, capturing intricate geometric relationships necessitates scaling network dimensions, leading to parameter explosion and scaling challenges.
@@
In contrast, the recently proposed Kolmogorov-Arnold Network (KAN)  (Liu et al. 2024) offers a promising MLP alternative, gaining significant attention for its accuracy and parameter efficiency. KAN leverages the Kolmogorov-Arnold Representation Theorem (KART)  (Arnold 2009), which states that any multivariate continuous function can be decomposed into univariate functions. KAN replaces fixed activation functions with learnable univariate functions (e.g., B-splines), directly modeling non-linearities and reducing reliance on redundant linear layers. This makes KAN particularly suitable for capturing hierarchical geometric features in irregular data like point clouds.
@@
Building on this foundation, we propose PointKAN, a KAN-based architecture for point cloud analysis. Point clouds are mapped to high-dimensional space and processed by a Geometric Affine Module (GAM). GAM utilizes Group-Norm for point grouping, normalization, and affine transformation, alongside S-Pool (employing Softmax-Pooling) to aggregate group-wise global information. These parallel outputs feed into KAN layers to extract complementary local and global features. Local features undergo max-pooling and are fused with global information, followed by a Residual Point (ResP) Block for deep feature aggregation. To address KAN’s computational inefficiency and parameter growth, PointKAN-elite integrates Efficient-KAN, inspired by Group-Rational KANs  (Yang and Wang 2024). Trained from scratch, PointKAN achieves excellent performance across point cloud datasets (Fig 1(a)), surpassing even pre-trained Point-BERT  (Yu et al. 2022) in few-shot learning (Fig 1(b)). It also demonstrates significant reductions in parameters and FLOPs (Fig 1(c-d)), highlighting KAN’s potential for 3D point cloud processing.
@@ 2 Related Work
2.1 A Review of PointMLP
Traditional point cloud analysis methods (e.g., graph convolution  (Wang et al. 2019) or attention mechanisms  (Zhao et al. 2021)) involve complex local geometric operations, resulting in high computational costs and limited generalization capability.
PointMLP  (Ma et al. 2022) simplifies network architecture to enhance efficiency and performance. It employs a multi-stage design where Farthest Point Sampling (FPS) and K-Nearest Neighbors (KNN) progressively reduce point density at each stage. Local features are extracted via Residual MLP Blocks—comprising linear layers, normalization, and activations with residual connections to mitigate gradient vanishing. Stacking these stages forms a deep hierarchical network for point cloud analysis. The core operation is formulated as:
$g_{i}=\Phi_{pos}(\mathcal{A}(\Phi_{pre}(f_{i,j}),|j=1,\cdots,K)),$ (1)
where $\Phi_{pre}$ and $\Phi_{pos}$ are residual MLP blocks, $\mathcal{A}$ denotes max-pooling and $f_{i,j}$ represents the feature of the $j$-th point in the $i$-th group, with each group encompassing $K$ points. Despite its simple architecture, PointMLP (Ma et al. 2022) delivers excellent performance on benchmarks like ModelNet40 (Wu et al. 2015) and ScanObjectNN (Uy et al. 2019).
@@ 2.2 Kolmogorov-Arnold Network
Kolmogorov-Arnold Networks (KAN)  (Liu et al. 2024) represent a novel architecture grounded in the Kolmogorov-Arnold Representation Theorem (KART)  (Arnold 2009). Unlike MLP, KAN replaces fixed activation functions with learnable univariate functions. KART establishes that any multivariate continuous function decomposes into finite univariate functions. KAN implements this using B-spline bases and compositional operations, providing a flexible and interpretable framework for high-dimensional function approximation.
KAN demonstrates growing efficacy across domains including computer vision  (Ferdaus et al. 2024; Ma et al. 2025; Moradi et al. 2024; Mohan, Wang, and Zhu 2024) and medical imaging  (Yang et al. 2025; Agrawal et al. 2024). In vision, KAN-augmented convolutional layers replace linear transformations with learnable per-pixel nonlinear activations  (Bodner et al. 2024), enhancing parameter richness and expressiveness for image recognition. Medical implementations integrate KAN layers into frameworks like U-Net  (Ronneberger, Fischer, and Brox 2015), improving segmentation accuracy and interpretability  (Wu et al. 2024; Li et al. 2024; Tang, Chen, and Shu 2024). For point clouds, PointNet-KAN  (Kashefi 2024) shows competitive classification and segmentation results versus PointNet  (Qi et al. 2017a). However, PointNet’s inherent limitations constrain this benchmark, and KAN’s potential for complex point cloud tasks remains largely unexplored. Effective integration of KAN with advanced point cloud architectures thus presents a critical research direction.
@@ 3 Method
We propose leveraging KAN for joint local-global feature learning in point clouds. To achieve this, we first review the Kolmogorov-Arnold Representation Theorem (KART) and KAN fundamentals(section 3.1), then detail our model’s key designs(section 3.2) including Efficient-KAN(section 3.3) in PointKAN-elite.
@@ 3.1 Preliminaries: KART and KAN
where $\psi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ are univariate continuous functions. Although eq. 2 theoretically enables high-dimensional function representation via two univariate layers, these functions may be potentially highly complex (even fractal) and thus unlearnable by neural networks.
KAN’s innovation overcomes KART’s two-layer function limitation by implementing a multi-layer network trained via backpropagation. A continuous and smooth univariate spline function $spline(x)$ is generated by the linear combination of B-spline base functions ($B_{i}(x)$), as shown in eq. 3. The coefficients $c_{i}$ therein are learnable. The activation function $\phi(x)$ is the sum of $silu(x)$ and $spline(x)$, where $silu(x)$ acts similar to a residual connection, as shown in eq. 4.
$spline(x)=\sum_{i}c_{i}B_{i}(x),$ (3)
$\phi(x)=silu(x)+spline(x),$ (4)
@@ 3.2 PointKAN
Despite KAN’s superior high-dimensional approximation capabilities and parameter efficiency, its adaptation to 3D point clouds remains challenging. The discrete spline nodes in KAN’s activation functions struggle to comprehensively capture local geometric features, while per-dimension parameter storage incurs significant memory overhead for large-scale networks. Additionally, B-spline computations exhibit suboptimal hardware acceleration, leading to inference latency. To overcome these limitations, we propose PointKAN—featuring a geometric affine module and parallel local feature extraction structure—alongside its lightweight variant PointKAN-elite, which jointly reduce memory footprint and accelerate training and inference speed. The complete pipeline is illustrated in fig. 2.
Geometric Affine Module
This module integrates Group-Norm and S-Pool for enhanced local feature extraction. Given input point cloud data $P\in\mathbb{R}^{N\times 3}$, we first apply Farthest Point Sampling (FPS) to select $G$ center points $P_{C}\in\mathbb{R}^{G\times 3}$, then construct local groups $\{x^{i}_{p}\in\mathbb{R}^{K\times 3}\}_{i=1}^{G}$ around each center using K-Nearest Neighbors (KNN). A lightweight PointNet (Qi et al. 2017a) processes each group to generate features $\{f^{i}_{j}\}_{j=1}^{K}\in\mathbb{R}^{K\times d}$ (where $d$ denotes feature dimension) and center features $\{f^{i}\}_{i=1}^{G}\in\mathbb{R}^{G\times d}$, which collectively form the Group-Norm input.
To enhance model stability and robustness, we design the Group-Norm module performing intra-group feature normalization, learnable affine transformation, and central-point feature propagation. Features $\{f^{i}_{j}\}_{j=1}^{K}$ are normalized:
@@
To address the lack of global context from mutually independent group features when directly feeding Group-Norm outputs to KAN, we introduce S-Pool in the GAM. This aggregates group-wise features, with S-Pool and Group-Norm outputs processed in parallel to supplement global information. Unlike max-pooling which maintains permutation invariance but incurs information loss, our softmax-inspired S-Pool maximizes intra-group feature preservation as defined in eq. 6,
@@
Local Feature Processing
In the Local Feature Processing (LFP), the two outputs of the GAM are processed in parallel using the KAN block. The KAN block, containing the KAN layer, effective depth-wise convolution (Chollet 2017) layer (DwConv), and residual connection, adds DwConv after the KAN layer to help KAN learn rich feature representations from high-dimensional channel information. Let $\mathcal{F}(\cdot)$ represent the processing of the KAN block, which can be denoted as:
@@
where ${f_{in}}$ denotes the input point cloud features, $\Phi$ represents the KAN layer. The shared LFP module is designed to learn local features from grouped point clouds. After max-pooling, it is added to the global information learned from the central point clouds of each group. This makes the aggregated feature information outputted by each group of point clouds more abundant.
Global Feature Processing
The Global Feature Processing (GFP) module utilizes Residual Point (ResP) Blocks to extract deep aggregated features, leveraging MLP’s inherent capacity for cross-channel fusion through fully-connected layers—a capability where KAN exhibits limitations despite their strength in capturing high-frequency details via univariate function decomposition. The ResP Block’s structure $x+MLP(x)$, with MLP comprising two sequential stacks of linear layers, normalization, and activations, provides critical advantages: its permutation invariance satisfies point cloud processing requirements, residual connections enable deep feature extraction, and purely feed-forward operations ensure computational efficiency for massive data. These properties facilitate efficient stacking of multiple ResP Blocks within GFP, as validated by ablation studies (see supplementary materials).

### v4

@@ Abstract
Multi-Layer Perceptron (MLP) has become one of the fundamental architectural component in point cloud analysis due to its simple, flexible structure and universality. However, when processing complex geometric structures in point cloud, MLP’s fixed activation functions struggle to efficiently capture local geometric features, while suffering from poor parameter and computational efficiency. The high redundancy in deep MLP structures imposes constraints on the model’s generalizability. In this paper, we propose PointKAN, which applies Kolmogorov-Arnold Network (KAN) to point cloud analysis tasks to investigate their efficacy in hierarchical feature representation. PointKAN adopts a hierarchical structure that progressively expands the receptive field, with each layer comprising three components: Geometric Affine Module (GAM), Local Feature Processing (LFP), and Global Feature Processing (GFP). The LFP stage employs a KAN structure to enhance feature extraction capabilities. Experimental results demonstrate that PointKAN outperforms PointMLP on benchmark datasets such as ModelNet40, ScanObjectNN, and ShapeNetPart. Notably, when handling large-scale point cloud data, PointKAN achieves significantly reduced computational overhead compared to PointMLP. Furthermore, its outstanding performance in few-shot learning tasks indicates that PointKAN exhibits stronger generalization capabilities than PointMLP. To further enhance parameter and computational efficiency, we develop Efficient-KAN in the PointKAN-elite variant. This work highlights the unique advantages of KAN over MLP in point cloud analysis and opens new avenues for research in point cloud understanding. Our code is available at https://github.com/Shiyan-cps/PointKAN-pytorch.
Figure 1: PointKAN (including PointKAN-elite) comprehensively outperforms counterparts. (a) Surpasses MLP/CNN baselines across datasets; (b) In Few-shot Learning, exceeds Point-BERT’s pre-trained performance when trained from scratch; (c)-(d) Substantially reduces parameters versus PointMLP, with PointKAN-elite significantly curbing FLOPs growth as Patch Sequence Length increases.
@@ 1 Introduction
Point cloud analysis is vital in computer vision and 3D processing, with applications spanning autonomous driving  (Li et al., 2023) and robotics (Yan et al., 2020; Rusu & Cousins, 2011). Unlike structured 2D images, point clouds pose challenges due to their irregularity and sparsity. Recent deep learning advances (Zhang et al., 2022; Pang et al., 2022; He et al., 2024; Liang et al., 2024; Han et al., 2024) build upon pioneering MLP-based architectures like PointNet/PointNet++  (Qi et al., 2017a; b). Subsequent research, including PointNeXt  (Qian et al., 2022), DGCNN  (Wang et al., 2019), and Point Transformer  (Zhao et al., 2021), employs CNN, GCN, and attention to capture local geometric features.
@@
Notably, PointMLP (Ma et al., 2022) utilizes a pure MLP architecture for point cloud analysis. Its core design philosophy eschews complex convolutions or graph operations, instead relying on stacked residual MLP blocks for feature learning. While achieving competitive performance in classification and segmentation, PointMLP inherits inherent MLP limitations. Rooted in the Universal Approximation Theorem (Cybenko, 1989), MLPs employ linear weights and fixed activations. Consequently, capturing the geometric relationships in large-scale point cloud data requires an increase in the number of hidden layers, which inevitably leads to reduced computational efficiency and model redundancy.
@@
In contrast, the recently proposed Kolmogorov-Arnold Network (KAN)  (Liu et al., 2024) offers a promising MLP alternative, gaining significant attention for its accuracy and interpretability. KAN leverages the Kolmogorov-Arnold Representation Theorem (KART)  (Arnold, 2009), which states that any multivariate continuous function can be decomposed into univariate functions. KAN replaces fixed activation functions with learnable univariate functions (e.g., B-splines), directly modeling non-linearities and reducing reliance on redundant linear layers. This makes KAN particularly suitable for capturing hierarchical geometric features in irregular data like point clouds. In terms of parameter efficiency, KAN requires more parameters than MLP for models of the same scale. However, fortunately, smaller-scale KANs exhibit better generalization capabilities. The parameters in KAN are utilized to shape the activation functions, rather than relying on increasing the number of neurons, as in MLP, to enhance model capacity.
@@
Building on this foundation, we propose PointKAN, a point cloud analysis framework that leverages the Kolmogorov-Arnold Network (KAN). Point clouds are first mapped to a high-dimensional space and preprocessed by the Geometric Affine Module (GAM), which performs affine transformations and Softmax pooling (S-Pool) operations. Local features are extracted in parallel by the Local Feature Processing Module, and the aggregated information is passed to the Residual Point (ResP) module to extract global features. To further enhance the computational efficiency of KAN and reduce the number of parameters, the PointKAN-elite version incorporates Efficient-KAN, inspired by Group-Rational KANs (Yang & Wang, 2024). Trained from scratch, PointKAN achieves excellent performance across point cloud datasets (Figure 1(a)), surpassing even pre-trained Point-BERT  (Yu et al., 2022) in few-shot learning (Figure 1(b)). It also demonstrates significant reductions in parameters and FLOPs (Figure 1(c-d)), highlighting KAN’s potential for 3D point cloud processing.
@@ 2 Related Work
2.1 Point Cloud Analysis
The analysis of point clouds has evolved from handcrafted features to deep learning. PointNet (Qi et al., 2017a) pioneered direct processing of unordered points, while PointNet++ (Qi et al., 2017b) introduced hierarchical feature learning. This progress spawned three major paradigms: point-based convolution (Li et al., 2018; Thomas et al., 2019), graph convolution (Wang et al., 2019), and sparse voxelization (Maturana & Scherer, 2015). In recent years, Transformer architectures (Zhao et al., 2021; Guo et al., 2021) have been incorporated to effectively capture long-range dependencies. These models often involve complex local geometric operations that result in high computational overhead, thereby limiting their practical applicability.
PointMLP  (Ma et al., 2022) simplifies network architecture to enhance efficiency and performance. It employs a multi-stage design where Farthest Point Sampling (FPS) and K-Nearest Neighbors (KNN) progressively reduce point density at each stage. Local features are extracted via Residual MLP Blocks—comprising linear layers, normalization, and activations with residual connections to mitigate gradient vanishing. Stacking these stages forms a deep hierarchical network for point cloud analysis. The core operation is formulated as $g_{i}=\Phi_{pos}(\mathcal{A}(\Phi_{pre}(f_{i,j}),|j=1,\cdots,K))$, where $\Phi_{pre}$ and $\Phi_{pos}$ are residual MLP blocks, $\mathcal{A}$ denotes max-pooling and $f_{i,j}$ represents the feature of the $j$-th point in the $i$-th group, with each group encompassing $K$ points. Despite its simple architecture, PointMLP (Ma et al., 2022) delivers excellent performance on benchmarks like ModelNet40 (Wu et al., 2015) and ScanObjectNN (Uy et al., 2019).
@@ 2.2 Kolmogorov-Arnold Network
Kolmogorov-Arnold Network (KAN)  (Liu et al., 2024) represent a novel architecture grounded in the Kolmogorov-Arnold Representation Theorem (KART)  (Arnold, 2009). Unlike MLP, KAN replaces fixed activation functions with learnable univariate functions. KART establishes that any multivariate continuous function decomposes into finite univariate functions. KAN implements this using B-spline bases and compositional operations, providing a flexible and interpretable framework for high-dimensional function approximation.
KAN demonstrates growing efficacy across domains including computer vision  (Ferdaus et al., 2024; Ma et al., 2025; Moradi et al., 2024; Mohan et al., 2024) and medical imaging  (Yang et al., 2025; Agrawal et al., 2024). In vision, KAN-augmented convolutional layers replace linear transformations with learnable per-pixel nonlinear activations  (Bodner et al., 2024), enhancing accuracy and expressiveness for image recognition. Medical implementations integrate KAN layers into frameworks like U-Net  (Ronneberger et al., 2015), improving segmentation accuracy and interpretability  (Wu et al., 2024; Li et al., 2024; Tang et al., 2024). For point clouds, PointNet-KAN  (Kashefi, 2024) shows better classification and segmentation results versus PointNet  (Qi et al., 2017a). However, PointNet’s inherent limitations constrain this benchmark, and KAN’s potential for complex point cloud tasks remains largely unexplored. Effective integration of KAN with advanced point cloud architectures thus presents a critical research direction.
@@ 3 Method
We propose leveraging KAN for local feature learning in point clouds. To achieve this, we first review the Kolmogorov-Arnold Representation Theorem (KART) and KAN fundamentals(Section 3.1), then detail our model’s key designs(Section 3.2) including Efficient-KAN(Section 3.3) in PointKAN-elite.
@@ 3.1 Preliminaries: KART and KAN
where $\psi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ are univariate continuous functions. Although Equation 1 theoretically enables high-dimensional function representation via two univariate layers, these functions may be highly complex (even fractal) and thus unlearnable by neural networks.
KAN’s innovation overcomes KART’s two-layer function limitation by implementing a multi-layer network trained via backpropagation. A continuous and smooth univariate spline function $spline(x)$ is generated by the linear combination of B-spline base functions ($B_{i}(x)$), as shown in Equation 2. The coefficients $c_{i}$ therein are learnable. The activation function $\phi(x)$ is the sum of $silu(x)$ and $spline(x)$, where $silu(x)$ acts similar to a residual connection.
$spline(x)=\sum_{i}c_{i}B_{i}(x),\quad\ \phi(x)=silu(x)+spline(x),$ (2)
@@ 3.2 PointKAN
Despite KAN’s superior high-dimensional function approximation capabilities and parameter efficiency, its adaptation to 3D point clouds remains challenging. The activation functions in KAN are obtained through a linear combination of basis functions, making them sensitive to input variations and consequently reducing the robustness of the model, while per-dimension parameter storage incurs significant memory overhead for large-scale networks. Additionally, B-spline computations exhibit suboptimal hardware acceleration, leading to inference latency. To overcome these limitations, we propose PointKAN—featuring a geometric affine module and parallel local feature extraction structure—alongside its lightweight variant PointKAN-elite, which reduces memory footprint and accelerate training and inference speed. The complete pipeline is illustrated in Figure 2.
3.2.1 Geometric Affine Module
This module integrates Group-Norm and S-Pool for enhanced model robustness. Given input point cloud data $P\in\mathbb{R}^{N\times 3}$, we first apply Farthest Point Sampling (FPS) to select $G$ center points $P_{C}\in\mathbb{R}^{G\times 3}$, then construct local groups $\{x^{i}_{p}\in\mathbb{R}^{K\times 3}\}_{i=1}^{G}$ around each center using K-Nearest Neighbors (KNN). A lightweight PointNet (Qi et al., 2017a) processes each group to generate features $\{f^{i}_{j}\}_{j=1}^{K}\in\mathbb{R}^{K\times d}$ (where $d$ denotes feature dimension) and center features $\{f^{i}\}_{i=1}^{G}\in\mathbb{R}^{G\times d}$, which collectively form the Group-Norm input.
Group-Norm module performs intra-group feature normalization, learnable affine transformation, and central-point feature propagation. Features $\{f^{i}_{j}\}_{j=1}^{K}$ are normalized:
@@ 3.2.1 Geometric Affine Module
To address the lack of global context from mutually independent group features when directly feeding Group-Norm outputs to KAN, we introduce S-Pool in the GAM. This aggregates group-wise features, with S-Pool and Group-Norm outputs processed in parallel to supplement global information. Unlike max-pooling which maintains permutation invariance but incurs information loss, our softmax-inspired S-Pool maximizes intra-group feature preservation as defined in Equation 4,
@@
3.2.2 Local Feature Processing
Compared to MLP, KAN employs learnable activation functions, as shown in Equation 2. The key feature of this structure lies in the fact that each spline $B_{i}(x)$ has a local support region. For a given input $x$ , only the basis functions $B_{i}(x)$ near $x$ are explicitly activated (with larger $c_{i}$ values), while the contributions of $B_{i}(x)$ from distant regions can be ignored. This characteristic allows KAN to exhibit excellent sensitivity to local variations and enables adjustments through hyperparameters such as kk(basis function order) and GG(grid size) to achieve an appropriate granularity of fitting. The local support property ensures that the model remains focused on the regions around the input $x$, making KAN suitable for structured geometric representations.
Figure 3: Local geometric feature sensitivity comparison. KAN exhibits superior sensitivity to local geometric variations in point clouds compared to MLP.
To validate that KAN is more suitable for local feature extraction than MLP, we feed point clouds into LFP modules with different architectures and compute the summation of the L2-norm for the output features. Points with values greater than a given threshold are highlighted in red. As depicted in Figure 3, we can intuitively observe that KAN exhibits greater sensitivity to edge points, abrupt change points, and junction points between different parts of the point cloud, demonstrating its superior capability in learning local feature variations.
The LFP structure is depicted in Figure 2(c). Let $\mathcal{F}(\cdot)$ represent the processing procedure of the KAN block, which is expressed as:
@@ 3.2.2 Local Feature Processing
where $f_{in}$} denotes the input point cloud features, $\Phi$ represents the KAN layer , and DwConv is a depthwise convolution, which assists KAN in learning rich feature representations from high-dimensional information.
3.2.3 Global Feature Processing
The GFP structure, as shown in Figure 2(e), fully leverages the ability of MLP to achieve cross-channel fusion within fully connected layers—MLP focuses on reweighting features across channels during aggregation, rather than directly adding the results of activation function processing as in KAN.(More detailed explanations are provided in the appendix) This structure satisfies the invariance requirements of point cloud processing, and the purely feed-forward operations ensure computational efficiency in large-scale data processing.